A. Two oncologists, Drs. Nodule and Smith independently examined 200 biopsies for evidence of invasive breast cancer. Dr. Nodule diagnosed 20 biopsies as “positive” for invasive breast cancer. Dr. Smith diagnosed 15 biopsies as “positive” for invasive breast cancer, 10 of which were also diagnosed as positive by Dr. Nodule. Calculate the % positive agreement between Drs. Nodule and Smith. Please round to the nearest percent. a. 93% b. 50% c. 43% d. 40% e. 34%

2.The incidence of heart attacks among smokers in a community is 30.0 per 100,000 per year. The incidence of heart attacks among nonsmokers in this community is 4.0 per 100,000 per year. The proportion of nonsmokers in this community is 75%. What proportion of the risk of heart attacks among smokers can be attributed to smoking? (Select one) a. 6.5% b. 26.0% c. 76.0% d. 86.7%

3.What proportion of the risk of heart attacks in the total population can be attributed to smoking? (Select one) a. 21.3% b. 61.9% c. 82.2% d. Cannot be calculated given the information provided.
Answer #1

1. Percentage agreement:

(agreements/(agreements + disagreements)) x 100% is the percentage agreement. Drs. Nodule and Smith have an agreement of invasive breast cancer in 10 biopsies. They have a disagreement on 10 biopsies of Drs. Nodule and 5 biopsies of Dr. Smith. So, the disagreement is for 15 biopsies. Hence,

% positive agreement between Drs. Nodule and Smith = (10/(10+15))*100=(10/20)*100

=40%

Ans: d. 40%